OpenStudy (anonymous):

sin ( a+ b)- sin (a-b)= 2 cos a sin B show work. help me please

jimthompson5910 (jim_thompson5910):

hints: sin(a+b) = sin(a)cos(b) + cos(a)sin(b) sin(a-b) = sin(a)cos(b) - cos(a)sin(b)

OpenStudy (anonymous):

thank you :)

jimthompson5910 (jim_thompson5910):

you're welcome